# Traces on the uniform tracial completion of $\mathcal{Z}$-stable C-algebras

### v2

@@ 1. Introduction
When $A$ is amenable, the von Neumann algebra $\pi_{\tau}(A)^{\prime\prime}$ is hyperfinite by Connes’ theorem ([14]). A breakthrough of Matui and Sato established lifting techniques for deducing structural results about a simple nuclear C-algebra $A$ from properties of the von Neumann algebras $\pi_{\tau}(A)^{\prime\prime}$ when $T(A)$ was non-empty and finite-dimensional ([33]). These ideas were extended by several authors to certain infinite-dimensional trace spaces ([46, 32, 42]). The main difficulty is that now one needs to work with the uniform 2-norm $\|a\|_{2,T(A)}=\sup_{\tau\in T(A)}\|a\|_{2,\tau}$ and the infinitely many extreme points of $T(A)$ will no longer be a topologically discrete set.
@@
The Jiang–Su algebra $\mathcal{Z}$ ([30]), which appears in Theorem A, plays a fundamental role in the classification of simple nuclear C-algebras since both $A$ and $A\otimes\mathcal{Z}$ have the same K-theory and trace space. Tensorial absorption of $\mathcal{Z}$, also known as $\mathcal{Z}$-stability, has emerged as the key regularity property in the Elliott classification programme ([22]).
The trace problem was previously known to have a positive solution for finite-dimensional trace simplices, where $\overline{A}{}^{T(A)}$ is a finite direct sum of finite factors and so von Neumann algebraic methods suffice. Beyond this point, $\overline{A}{}^{T(A)}$ is no longer a von Neumann algebra, so a solution to the trace problem has to take the topology of $T(A)$ and the bundle-like structure of $\overline{A}{}^{T(A)}$ into account. Indeed, trace problem was open even for approximately finite-dimensional algebras with a compact (but infinite) set of extreme traces.
@@
Beyond regularity, none of the other assumptions typically seen in the Elliott classification programme are needed to establish Theorem A. In particular, it holds for non-simple C-algebras and non-nuclear C-algebras. Moreover, $\mathcal{Z}$-stability only enters in order to ensure the existence of complemented partitions of unity ([12, 9]), which we will discuss below, and to rule out type I quotients. In particular, uniform property $\Gamma$ would be a suitable alternative regularity property ([11]).
@@ Acknowledgements
I’d like to thank Aaron Tikuisis, Hannes Thiel, Andrea Vaccaro and Stuart White for their comments on an earlier version of this manuscript.
@@ 2.1. Traces and Choquet simplices
A Choquet simplex $X$ is a compact convex set where every $x\in X$ is the barycentre of a unique Radon probability measure $\mu_{x}$ supported on the extreme boundary $\partial_{e}X$ of $X$; see for example [1, Theorem II.3.6]. [^1] When $X$ is non-metrizable, $\partial_{e}X$ need not be Borel, so supported on the extreme boundary should be interpreted as $\mu_{x}(E)=0$ for any Borel measurable set $E$ that doesn’t intersect $\partial_{e}X$. The trace space $T(A)$ of a unital C-algebra is a Choquet simplex by [41, Theorem 3.1.18]. When $A$ is non-unital, $T(A)$ need not be compact; however, if $T(A)$ is compact, then $T(A)$ is a Choquet simplex; see for example [8, Theorem 2.6].
In this subsection, we recall some important facts about continuous affine function on Choquet simplices in general and on $T(A)$ in particular. We write $\mathrm{Aff}(X)$ for the space of continuous affine functions $X\rightarrow\mathbb{R}$, $\mathrm{Aff}(X)_{+}$ for the non-negative valued affine functions $X\rightarrow[0,\infty)$, and $\|\cdot\|_{\infty}$ for the supremum norm on $\mathrm{Aff}(X)$. The following result is well-known and is often attributed to Cuntz and Pedersen ([17]); see [32, Lemma 6.2] or [7, Proposition 2.1] for a proof.
@@ 2.3. Tracially complete C-algebras
In this paper, we work in the general framework of tracially complete C-algebras, introduced in [8]. This framework includes as special cases both the uniform tracial completions of C-algebras discussed above and Ozawa’s W-bundles, introduced in [38].
@@ 2.4. Finite von Neumann algebras
Lemma 2.7 ([44, Chapter XIV, Lemma 2.2]).
@@ 2.5. Strict comparison
for all $a\in(A\otimes\mathbb{K})_{+}$ and $\epsilon>0$. Endowed with this topology, $\mathrm{QT}_{c}(A)$ is a compact; see [21, Section 4.1]. The subspace topology on $T(A)$ inherited from $\mathrm{QT}_{c}(A)$ and the subspace topology on $T(A)$ induced by the weak-topology on $A^{*}$ coincide; see [21, Proposition 3.10]. [^2] We warn the reader that [21] uses a different notational convention: $T(A)$ denotes cone of extended traces and the bounded traces are $T_{I}(A)$ where $I=A$. In particular, $T(A)$ is a compact subset of $\mathrm{QT}_{c}(A)$ when $A$ is unital.
@@
for all $\tau\in K$. Choose $\lambda\in\Lambda$ such that $\|(a-\epsilon)_{+}-e_{\lambda}(a-\epsilon)_{+}e_{\lambda}\|<\epsilon$ and $\|b-e_{\lambda}be_{\lambda}\|<\delta$. Then, by [40, Proposition 2.2], we have
@@ 3. The main construction
The conclusion of Theorem 3.7 could be interpreted informally as saying that $(\mathcal{M},X)$ has real rank zero at the level of the Cuntz semigroup. Indeed, if $A$ is a C-algebra with real rank zero and $a\in A_{+}$, then the hereditary subalgebra $\overline{aAa}$ has an approximate unit consisting of projections, so there exists a projection $p\in A$ with $(a-\epsilon)_{+}\precsim p\precsim a$ by [40, Proposition 2.2]. Passing to the stabilisation, this argument shows that the Cuntz semigroup $\mathrm{Cu}(A)$ of a C-algebra with real rank zero is algebraic in the sense of [3, Section 5.5], i.e. every element of $\mathrm{Cu}(A)$ is the supremum of a sequence of compact elements. This observation goes back to [15, Corollary 5], where it is also shown that the converse holds when $A$ has stable rank one.
@@ 4. The main theorems
Since $\mathcal{M}$ has strict comparison with respect to $X$, it also has strict comparison with respect to the (larger) set of all lower semi-continuous extended traces on $\mathcal{M}$. Hence, all lower semi-continuous 2-quasitraces on $\mathcal{M}$ are additive by [37, Theorem 3.6(i)]. ∎

### v3

@@ 1. Introduction
When $A$ is amenable, the von Neumann algebra $\pi_{\tau}(A)^{\prime\prime}$ is hyperfinite by Connes’ theorem ([14]). A breakthrough of Matui and Sato established lifting techniques for deducing structural results about a simple nuclear C-algebra $A$ from properties of the von Neumann algebras $\pi_{\tau}(A)^{\prime\prime}$ when $T(A)$ was non-empty and finite-dimensional ([34]). These ideas were subsequently extended to cover certain infinite-dimensional trace spaces ([48, 33, 43]) and, when $A$ also has stable rank one and locally finite nuclear dimension, to all trace spaces ([46]).
@@
The Jiang–Su algebra $\mathcal{Z}$ ([31]), which appears in Theorem A, plays a fundamental role in the classification of simple nuclear C-algebras since both $A$ and $A\otimes\mathcal{Z}$ have the same K-theory and trace space. Tensorial absorption of $\mathcal{Z}$, also known as $\mathcal{Z}$-stability, has emerged as the key regularity property in the Elliott classification programme ([22]). In particular, all simple approximately finite-dimensional C-algebras are $\mathcal{Z}$-stable.
The condition that $T(A)$ is compact is automatic when $A$ is unital. Theorem A is formulated so that it also covers a wide class of non-unital examples. In the non-unital case, the trace space $T(A)$ is defined as the space of positive tracial functionals $\tau\in A^{*}$ with $\|\tau\|_{A^{*}}=1$ and it may or may not be compact. However, every simple separable exact $\mathcal{Z}$-stable C-algebra is stably isomorphic to one with compact trace space (see for example [9, Theorem 2.7]).
The trace problem was previously known to have a positive solution for finite-dimensional trace simplices, where $\overline{A}{}^{T(A)}$ is a finite direct sum of finite factors and so von Neumann algebraic methods suffice. Beyond this point, $\overline{A}{}^{T(A)}$ is no longer a von Neumann algebra, so a solution to the trace problem has to take the topology of $T(A)$ and the bundle-like structure of $\overline{A}{}^{T(A)}$ into account. Indeed, the trace problem was open even for approximately finite-dimensional algebras with a compact (but infinite) set of extreme traces.
@@
Beyond regularity, none of the other assumptions typically seen in the Elliott classification programme are needed to establish Theorem A. In particular, it holds for non-simple C-algebras and non-nuclear C-algebras. Therefore, its future applications are not limited to the setting of the Elliott classification programme. Moreover, $\mathcal{Z}$-stability only enters in order to ensure the existence of complemented partitions of unity ([12, 9]), which we will discuss below, and to rule out type I quotients. In particular, uniform property $\Gamma$ would be a suitable alternative regularity property ([11]).
@@ Acknowledgements
I’d like to thank Aaron Tikuisis, Hannes Thiel, Andrea Vaccaro and Stuart White for their comments on an earlier version of this manuscript. I’d also like to thank the anonymous referee for their comments.
@@ 2.1. Traces and Choquet simplices
A Choquet simplex $X$ is a compact convex set where every $x\in X$ is the barycentre of a unique Radon probability measure $\mu_{x}$ supported on the extreme boundary $\partial_{e}X$ of $X$; see for example [1, Theorem II.3.6]. [^1] When $X$ is non-metrizable, $\partial_{e}X$ need not be Borel, so supported on the extreme boundary should be interpreted as $\mu_{x}(E)=0$ for any Baire measurable set $E$ that doesn’t intersect $\partial_{e}X$. The trace space $T(A)$ of a unital C-algebra is a Choquet simplex by [42, Theorem 3.1.18]. When $A$ is non-unital, $T(A)$ need not be compact; however, if $T(A)$ is compact, then $T(A)$ is a Choquet simplex; see for example [8, Theorem 2.6].
In this subsection, we recall some important facts about continuous affine function on Choquet simplices in general and on $T(A)$ in particular. We write $\mathrm{Aff}(X)$ for the space of continuous affine functions $X\rightarrow\mathbb{R}$, $\mathrm{Aff}(X)_{+}$ for the non-negative valued affine functions $X\rightarrow[0,\infty)$, and $\|\cdot\|_{\infty}$ for the supremum norm on $\mathrm{Aff}(X)$. The following result is well-known and is often attributed to Cuntz and Pedersen ([17]); see [7, Proposition 2.1] and [8, Proposition 2.7] for a proof.
@@ 2.3. Tracially complete C-algebras
In this paper, we work in the general framework of tracially complete C-algebras, introduced in [8] and further investigated in [26]. This framework includes as special cases both the uniform tracial completions of C-algebras discussed above and Ozawa’s W-bundles, introduced in [39].
@@ 2.4. Finite von Neumann algebras
Lemma 2.7 ([45, Chapter XIV, Lemma 2.2]).
@@ 2.5. Strict comparison
for all $a\in(A\otimes\mathbb{K})_{+}$ and $\epsilon>0$. Endowed with this topology, $\mathrm{QT}_{c}(A)$ is a compact Hausdorff cone; see [21, Section 4.1]. The subspace topology on $T(A)$ inherited from $\mathrm{QT}_{c}(A)$ and the subspace topology on $T(A)$ induced by the weak-topology on $A^{*}$ coincide; see [21, Proposition 3.10]. [^2] We warn the reader that [21] uses a different notational convention: $T(A)$ denotes the cone of extended traces and the bounded traces are $T_{I}(A)$ where $I=A$. In particular, $T(A)$ is a compact subset of $\mathrm{QT}_{c}(A)$ when $A$ is unital.
@@
for all $\tau\in K$. Choose $\lambda\in\Lambda$ such that $\|(a-\epsilon)_{+}-e_{\lambda}(a-\epsilon)_{+}e_{\lambda}\|<\epsilon$ and $\|b-e_{\lambda}be_{\lambda}\|<\delta$. Then, by [41, Proposition 2.2], we have
@@ 3. The main construction
The conclusion of Theorem 3.7 could be interpreted informally as saying that $(\mathcal{M},X)$ has real rank zero at the level of the Cuntz semigroup. Indeed, if $A$ is a C-algebra with real rank zero and $a\in A_{+}$, then the hereditary subalgebra $\overline{aAa}$ has an approximate unit consisting of projections, so there exists a projection $p\in A$ with $(a-\epsilon)_{+}\precsim p\precsim a$ by [41, Proposition 2.2]. Passing to the stabilisation, this argument shows that the Cuntz semigroup $\mathrm{Cu}(A)$ of a C-algebra with real rank zero is algebraic in the sense of [3, Section 5.5], i.e. every element of $\mathrm{Cu}(A)$ is the supremum of a sequence of compact elements. This observation goes back to [15, Corollary 5], where it is also shown that the converse holds when $A$ has stable rank one. In this language, Theorem 3.7 shows that $\mathrm{Cu}(\mathcal{M})$ is algebraic.
@@ 4. The main theorems
Since $\mathcal{M}$ has strict comparison with respect to $X$, it also has strict comparison with respect to the (larger) set of all lower semi-continuous extended traces on $\mathcal{M}$. Hence, all lower semi-continuous 2-quasitraces on $\mathcal{M}$ are additive by [38, Theorem 3.6(i)]. ∎